# Simulating and probing many-body quantum states in waveguide-QED systems with giant atoms

### v1

@@ Abstract
Waveguide quantum electrodynamics (wQED) with giant atoms provides a distinctive opportunity to study one-dimensional (1D) coupled spin systems through its unique decoherence-free interactions. This study presents a theoretical framework for simulating the diagonal Aubry-André-Harper (AAH) model in the context of giant-atom wQED. The proposed scheme employs photonic modes in the waveguide to not only mediate interactions between atoms but also to detect the energy spectrum of the atom array. To illustrate the effectiveness of this approach, we present a simulation of the Hofstadter butterfly spectrum with high precision. Furthermore, for an incommensurate AAH atomic chain, we demonstrate that the photon transmission spectrum can accurately distinguish between the many-body localized phase and the extended phase. The method presented here is also applicable to the simulation of other types of 1D atomic chains based on giant-atom wQED.
@@ I INTRODUCTION
Waveguide quantum electrodynamics (wQED) systems Roy et al. (2017); Gu et al. (2017); Sheremet et al. (2023) with giant atoms Kockum (2021) have emerged as a novel paradigm in quantum optics. In these systems, quantum systems are coupled to distant positions of the waveguide, which are separated by wavelength spacings. The structures of giant atoms have been realized in a variety of quantum systems, including superconducting qubits Andersson et al. (2019); Kannan et al. (2020) and giant spin ensembles Wang et al. (2022). Moreover, theoretical proposals have been put forth suggesting the use of cold atoms in optical lattices González-Tudela et al. (2019) and synthetic dimensions Du et al. (2022); Xiao et al. (2022) as a means of realizing the giant-atom configurations. It is evident that giant atoms cannot be regarded as pointlike particles, and thus the conventional dipole approximation is no longer applicable. The non-dipole effects of giant atom systems can produce a number of novel phenomena, including frequency-dependent decay rate and Lamb shift Kockum et al. (2014), non-Markovian dynamics Guo et al. (2017); Andersson et al. (2019); Guo et al. (2020a); Qiu et al. (2023); Lim et al. (2023); Xu and Guo (2024); Li and Shen (2024); Roccati and Cilluffo (2024), generation of enhanced entanglement Santos and Bachelard (2023); Yin and Liao (2023), tunable atom-photon bound states Guo et al. (2020b); Zhao and Wang (2020); Wang et al. (2021); Cheng et al. (2022); Zhang et al. (2023); Jia and Yu (2024) and scattering states Ask et al. ; Cai and Jia (2021); Feng and Jia (2021); Zhu et al. (2022); Yin et al. (2022); Peng and Jia (2023); Gu et al. (2023, 2024). One of the most unique properties of giant-atom wQED is the decoherence-free interaction between a pair of braided giant atoms coupling to a linear waveguide Kockum et al. (2018); Carollo et al. (2020), which has been successfully implemented in superconducting circuits Kannan et al. (2020). This type of protected exchange interactions between giant atoms can be used to perform coherent quantum operations Du et al. (2023); Soro et al. (2023). Another promising application of such decoherence-free interactions is quantum simulation of coupled spins, such as one-dimensional (1D) tight-binding atomic chain Kockum et al. (2018) and 1D Su-Schrieffer-Heeger (SSH) atomic chain Peng and Jia (2023).
@@
Quantum simulation provides a robust tool for investigating many-body phenomena that may present a significant challenge for classical computers Feynman (1982); Georgescu et al. (2014). The 1D Aubry-André-Harper (AAH) model Harper (1955); Aubry and André (1980), as a prominent tool for investigating topological physics and localization transition, has attracted considerable attention from theoretical perspectives Simon (1982); Thouless (1983); Ostlund et al. (1983); Hiramoto and Kohmoto (1989); Jitomirskaya (1999); Iyer et al. (2013); Ganeshan et al. (2013); Lellouch and Sanchez-Palencia (2014); Liu et al. (2015); Khemani et al. (2017). To illustrate, the diagonal AAH model can be precisely mapped to the two-dimensional (2D) Hofstadter model Hofstadter (1976), demonstrating a 2D quantum Hall effect (QHE) with topologically protected edge states, and the corresponding energy spectrum versus the dimensionless magnetic flux form a famous fractal structure known as the Hofstadter butterfly. In addition, the incommensurate diagonal AAH model describes a 1D tight-binding lattice with quasiperiodic potential, in which the disorder strength can induce a delocalization-localization phase transition Aubry and André (1980); Simon (1982); Jitomirskaya (1999). Experimentally, the AAH model was first implemented on quantum simulation platforms based on cold atoms in optical lattices Roati et al. (2008); Schreiber et al. (2015); Bordia et al. (2017) and photonic lattice Lahini et al. (2009); Kraus et al. (2012). More recently, with the development of modern nanotechnology, superconducting quantum circuits have proven to be an optimal platform for verifying the underlying physics of the AAH model Roushan et al. (2017); Shi et al. (2023); Li et al. (2023a, b).
@@
The inherent compatibility of giant-atom wQED systems and superconducting quantum circuits Andersson et al. (2019); Kannan et al. (2020) motivates us to investigate the feasibility of simulating AAH models based on giant-atom wQED. Therefore, in this paper, we construct the diagonal AAH model by employing the decoherence-free interactions that are unique to giant-atom wQED systems. In our proposed scheme, photonic modes in 1D waveguide serve not only to mediate interactions between atoms, but also to facilitate the readout of the spectral structure of atomic chain. By appropriately modulating the coupling configurations of the atomic chain, it is possible to reproduce the energy spectrum of the atomic chains with high precision through the photon scattering process. As an example, we demonstrate the efficacy of this method by simulating the Hofstadter butterfly spectrum at high resolution. Moreover, we find that for an incommensurate AAH atomic chain in the extended phase, the decays of the collective modes into the waveguide are concentrated in a few of them. In contrast, in the localized phase, the decays distribute evenly in each mode. This interesting correspondence making the localization transition easy to detect through the single-photon scattering spectrum. These results also indicate that wQED with giant atoms is optimal for investigating light-matter interactions between an AAH atomic chain and its surrounding photonic environment. The methodology presented here can be extended to simulating other types of 1D atomic chains.
@@
The remainder of this paper is organized as follows. In Sec. II, we derive the single-photon scattering amplitudes for wQED systems with multiple giant atoms based on the real-space formalism Shen and Fan (2005a, b). Additionally, we extract the effective Hamiltonian that describes the interacting atomic chain. In Sec. III, we discuss quantum simulation of AAH model through wQED systems with giant atoms. Specifically, we present a method for constructing an AAH-type atomic chain based on giant-atom wQED structure in Sec. III.1. Subsequently, we provide discussions on probing the energy band structure, simulating the Hofstadter butterfly, and probing the localization transition through single-photon scattering spectra in Sec. III.2-III.4. Additionally, we discuss the influence of spontaneous emission to the non-waveguide degrees of freedom in Sec. III.5. Finally, further discussions and conclusions are given in Sec. IV.
@@ II Theoretical description of single-photon scattering in waveguide QED containing multiple giant atoms
In this section, we provide an overview of the single-photon scattering problem for wQED with multiple giant atoms. From the derivation of the single-photon scattering amplitudes, we can extract the effective Hamiltonian describing the inter-atomic interactions mediated by the waveguide modes. This serves as the foundation for the construction of the target Hamiltonian that needs to be simulated. Additionally, in our proposal, single-photon scattering spectra are also employed as a means to probe the many-body states of atomic chains.
@@ II.1 Hamiltonian and equations of motion
Here we focus on the wQED structures comprising $N$ two-level giant atoms with modulated frequencies $\omega_{i}=\omega_{\mathrm{a}}+V_{0}\cos\left(2\pi\beta{i}+\varphi\right)$ ($i=1,2,\cdots,N$). $V_{0}$ is the modulation amplitude, $\beta$ controls the periodicity of the modulation, and $\varphi$ is the modulation phase. Each atom is coupled to a 1D waveguide through two connection points and each pair of neighboring atoms is in a braided configuration, as shown in Fig. 1(a). The connection points are located at position $x_{im}$ ($i=1,2,\cdots,N$, and $m=1,2$), the corresponding coupling strengths are $g_{i1}\equiv g_{1}$ and $g_{i2}\equiv g_{2}$, respectively. The phase delays between the coupling points are set as $\omega_{\mathrm{a}}(x_{i2}-x_{i1})/v_{\mathrm{g}}=\pi$ and $\omega_{\mathrm{a}}(x_{i2}-x_{i+1,1})/v_{\mathrm{g}}=\phi$ ($\phi\leq\pi/2$). $v_{\mathrm{g}}$ is the group velocity of the photons in the waveguide. Under the rotating-wave approximation, the Hamiltonian of the system in the real space can be written as ($\hbar=1$)
$\hat{H}=$ $\sum_{i=1}^{N}{\omega_{i}\hat{\sigma}_{i}^{+}\!\>\hat{\sigma}_{i}^{-}}+\int\mathrm{d}x\sum_{s}{\hat{c}_{s}^{\dagger}(x)\left(-\mathrm{i}l_{s}v_{\mathrm{g}}\frac{\partial}{\partial x}\right)\hat{c}_{s}\left(x\right)}$ (1)
$+\int\mathrm{d}x\sum_{s}\sum_{i=1}^{N}\sum_{m=1}^{2}{g_{m}}\delta\left(x-x_{im}\right)\left[\hat{c}_{s}^{\dagger}\left(x\right)\hat{\sigma}_{i}^{-}+\mathrm{H.c.}\right],$
where $s=\mathrm{R},\mathrm{L}$ and $l_{\mathrm{R},\mathrm{L}}=\pm 1$. $\hat{\sigma}_{i}^{+}$ ($\hat{\sigma}_{i}^{-}$) is the raising (lowering) operator of the atom $i$. $\hat{c}_{\mathrm{R}}^{\dagger}(x)$ [$\hat{c}_{\mathrm{R}}(x)$] and $\hat{c}_{\mathrm{L}}^{\dagger}(x)$ [$\hat{c}_{\mathrm{L}}(x)$] are the field operators of creating (annihilating) the right- and left-propagating photons at position $x$ in the waveguide.
@@
[…]
@@
$\left(-\mathrm{i}v_{\mathrm{g}}\frac{\partial}{\partial x}-\omega\right)\Phi_{\mathrm{R}}(x)+\sum_{i=1}^{N}\sum_{m=1}^{2}g_{m}\delta\left(x-x_{im}\right)f_{i}=0,$ (4a)
$\left(\mathrm{i}v_{\mathrm{g}}\frac{\partial}{\partial x}-\omega\right)\Phi_{\mathrm{L}}(x)+\sum_{i=1}^{N}\sum_{m=1}^{2}g_{m}\delta\left(x-x_{im}\right)f_{i}=0,$ (4b)
@@ II.2 Scattering amplitudes of photons and effective Hamiltonian of atom array
We assume that a single photon with energy $\omega=v_{\mathrm{g}}k$ is initially incident from the left, where $k$ is the wave vector of the photon. Then $\Phi_{\mathrm{R}}(x)$ and $\Phi_{\mathrm{L}}(x)$ take the following ansatz
$\Phi_{\mathrm{R}}(x)=e^{\mathrm{i}kx}\sum_{p=0}^{2N}{t_{p}}\vartheta\left(x-x_{p}\right)\vartheta\left(x_{p+1}-x\right),$ (5a)
$\Phi_{\mathrm{L}}(x)=e^{-\mathrm{i}kx}\sum_{p=1}^{2N}{r_{p}}\vartheta\left(x-x_{p-1}\right)\vartheta\left(x_{p}-x\right).$ (5b)
Here the positions of coupling points from left to right are labeled as $x_{1},x_{2},\cdots,x_{2N}$. Moreover, $x_{0}=-\infty$ and $x_{2N+1}=+\infty$ are defined. $t_{p}$ ($r_{p}$) is the transmission (reflection) amplitude of the $p$th coupling point. $t_{2N}\equiv t$ ($r_{1}\equiv r$) is the transmission (reflection) amplitude of the last (first) coupling point. $t_{0}=1$ is the amplitude of the incident field. $\vartheta(x)$ denotes the Heaviside step function.
Starting from the equations of motion (4a)-(4c) and the ansatz (5a)-(5b), and after some algebra Peng and Jia (2023), we can obtain the transmission and reflection amplitudes
@@
The transmittance and the reflectance can be further defined as $T=|t|^{2}$ and $R=|r|^{2}$, respectively. In Eqs. (6a) and (6b), $\Delta=\omega-\omega_{\mathrm{a}}$ is the photon-atom detuning, $\mathbf{I}$ is the identity matrix, and $\mathbf{V}$ takes the form
@@
Here the decay rate into the guided modes through the coupling point at $x_{im}$ is $\gamma_{m}=2g_{m}^{2}/v_{\mathrm{g}}$, and the phase factor is defined as $\theta_{im}=k_{\mathrm{a}}x_{im}=\omega_{\mathrm{a}}x_{im}/v_{\mathrm{g}}$. Note that we have assumed that the spacing between the connection points is small enough so that the phase-accumulated effects for detuned photons (the non-Markovian effects) can be neglected. Consequently, the wave vector $k$ in the definition of the phase factor has been replaced by $k_{\mathrm{a}}$. $\mathbf{H}$ is the effective non-Hermitian Hamilton matrix of the atom array, with elements
$\mathcal{H}_{ij}=\left(\omega_{i}-\omega_{\mathrm{a}}\right)\delta_{ij}-\frac{\mathrm{i}}{2}\sum_{m=1}^{2}\sum_{m^{\prime}=1}^{2}{{\sqrt{\gamma_{m}\gamma_{m^{\prime}}}}}e^{\mathrm{i}\left|\theta_{im}-\theta_{jm^{\prime}}\right|},$ (9)
which summarize the coherent and dissipative atom-atom interactions mediated by the waveguide modes. Moreover, this result indicates that the many-body model of the atomic chain can be designed by optimizing the layout of the connection points and the atom-waveguide coupling rates. Specifically, under parameters $\theta_{i2}-\theta_{i1}=\pi$ and $\theta_{i2}-\theta_{i+1,1}=\phi$, the effective Hamiltonian of the atom array reads
$\hat{H}_{\mathrm{eff}}$ $=$ $\sum_{i=1}^{N}\sum_{j=1}^{N}\mathcal{H}_{ij}\hat{\sigma}^{+}_{i}\hat{\sigma}^{-}_{j}$ (10)
$=$ $\sum_{i}\left[\left(\omega_{i}-\omega_{\mathrm{a}}\right)-\frac{\mathrm{i}}{2}\Gamma_{\mathrm{eff}}\right]\hat{\sigma}^{+}_{i}\hat{\sigma}^{-}_{i}$
$+\sum_{i\neq j}\left(G_{ij}-\frac{\mathrm{i}}{2}\Gamma^{(\mathrm{coll})}_{ij}\right)\hat{\sigma}^{+}_{i}\hat{\sigma}^{-}_{j}.$
Here the effective decay of each atom is defined as
$\Gamma_{\mathrm{eff}}=2\left(\gamma-\sqrt{\gamma^{2}-\delta^{2}}\right),$ (11)
@@
$G_{ij}=\left\{\begin{array}[]{ll}\gamma{\sin}\phi\equiv{J},&|i-j|=1\\ \frac{(-1)^{|i-j|-1}}{2}\Gamma_{\mathrm{eff}}\sin\left(|i-j|\phi\right),&|i-j|>1\end{array}\right.,$ (12a)
$\Gamma^{(\mathrm{coll})}_{ij}=(-1)^{|i-j|-1}\Gamma_{\mathrm{eff}}\cos[(i-j)\phi].$ (12b)
III Quantum simulation of AAH model through waveguide QED systems with giant atoms
Figure 2: (a) Tansmitance for an AAH giant-atom chain with $N=30$ sites as a function of detuning $\Delta$ and modulation phase $\varphi$ for different values of $\beta$. From left to right, $\beta$ is chosen as $1/2$, $1/3$, $1/4$, and $(\sqrt{5}-1)/2$, respectively. The phase delay is selected to be $\phi=\pi/2$, resulting in $J=\gamma$. Other parameters are set as $\delta=0.1\gamma$ and $V_{0}=2J$. (b) Energy spectra of an AAH model obtained from theoretical calculations for comparison with (a).
Figure 3: (a) Upper panel: transmission spectrum for an AAH giant-atom chain, with $\beta=1/4$ and $\varphi=3\pi/4$. Other parameters are the same as those used in Fig. 2 (a). Lower panel: the decomposed Lorentzian components for the spectrum shown in the upper panel. The resonances corresponding to the superradiant (subradiant) states in the band are labeled as SupS (SubS), with $\tilde{\Gamma}_{n}>\Gamma_{\mathrm{eff}}$ ($\tilde{\Gamma}_{n}<\Gamma_{\mathrm{eff}}$). The two edge states (ES) in the gaps exhibit a width that is approximately equal to $\Gamma_{\mathrm{eff}}$. (b) The center frequency $\tilde{\Delta}_{n}$ (the disks) of each Lorentzian component and the eigenenergy $E_{n}$ (the crosses) given by the exact AAH model. (c) The linewidth $\tilde{\Gamma}_{n}$ of each Lorentzian component. The effective decay $\Gamma_{\mathrm{eff}}$ of a single giant atom is labeled by the dashed line.

### v2

@@ Abstract
Waveguide quantum electrodynamics (wQED) with giant atoms provides a distinctive opportunity to quantum simulation of many-body physics through its unique decoherence-free interactions. This study presents a theoretical framework for simulating the diagonal Aubry-André-Harper (AAH) model in the context of giant-atom wQED. The proposed scheme employs photonic modes in the waveguide to not only mediate interactions between atoms but also to detect the energy spectrum of the atom array. To illustrate the effectiveness of this approach, we present a simulation of the Hofstadter butterfly spectrum with high precision. Furthermore, for an incommensurate AAH atomic chain, we demonstrate that the photon transmission spectrum can accurately distinguish between the many-body localized phase and the extended phase. The method presented here is also applicable to the simulation of other types of 1D atomic chains based on giant-atom wQED.
@@ I INTRODUCTION
Waveguide quantum electrodynamics (wQED) systems [1, 2, 3] with giant atoms [4] have emerged as a novel paradigm in quantum optics. In these systems, quantum systems are coupled to distant positions of the waveguide, which are separated by wavelength spacings. The structures of giant atoms have been realized in a variety of quantum systems, including superconducting qubits [5, 6] and giant spin ensembles [7]. Moreover, theoretical proposals have been put forth suggesting the use of cold atoms in optical lattices [8] and synthetic dimensions [9, 10] as a means of realizing the giant-atom configurations. It is evident that giant atoms cannot be regarded as pointlike particles, and thus the conventional dipole approximation is no longer applicable. The non-dipole effects of giant atom systems can produce a number of novel phenomena, including frequency-dependent decay rate and Lamb shift [11], non-Markovian dynamics [12, 5, 13, 14, 15, 16, 17, 18], generation of enhanced entanglement [19, 20], tunable atom-photon bound states [21, 22, 23, 24, 25, 26] and scattering states [27, 28, 29, 30, 31, 32, 33, 34]. One of the most unique properties of giant-atom wQED is the decoherence-free interaction between a pair of braided giant atoms coupling to a linear waveguide [35, 36], which has been successfully implemented in superconducting circuits [6]. This type of protected exchange interactions between giant atoms can be used to perform coherent quantum operations [37, 38]. Another promising application of such decoherence-free interactions is quantum simulation of coupled spins, such as one-dimensional (1D) tight-binding atomic chain [35], 1D Su-Schrieffer-Heeger atomic chain [32], and driven-dissipative spin chains [39].
@@
Quantum simulation provides a robust tool for investigating many-body phenomena that may present a significant challenge for classical computers [40, 41]. The 1D Aubry-André-Harper (AAH) model [42, 43], as a prominent tool for investigating topological physics and localization transition, has attracted considerable attention from theoretical perspectives [44, 45, 46, 47, 48, 49, 50, 51, 52, 53, 54]. To illustrate, the diagonal AAH model can be precisely mapped to the two-dimensional (2D) Hofstadter model [55], demonstrating a 2D quantum Hall effect (QHE) with topologically protected edge states. The corresponding energy spectra versus the dimensionless magnetic flux form a famous fractal structure known as the Hofstadter butterfly. In addition, the incommensurate diagonal AAH model describes a 1D tight-binding lattice with quasiperiodic potential, in which the disorder strength can induce a delocalization-localization phase transition [43, 44, 48]. Experimentally, the AAH model was first implemented on quantum simulation platforms based on cold atoms in optical lattices [56, 57, 58] and photonic lattice [59, 60]. More recently, with the development of modern nanotechnology, superconducting quantum circuits have proven to be an optimal platform for verifying the underlying physics of the AAH model [61, 62, 63, 64].
@@
The inherent compatibility between giant-atom wQED systems and superconducting quantum circuits [5, 6] motivates us to investigate the feasibility of simulating the AAH model based on giant-atom wQED. Therefore, in this paper, we construct the diagonal AAH model by employing the decoherence-free interactions that are unique to giant-atom wQED systems. In our proposed scheme, photonic modes in 1D waveguide serve not only to mediate interactions between atoms, but also to facilitate the readout of the spectral structure of atomic chain. By appropriately modulating the coupling configurations between the atoms and the waveguide, it is possible to reproduce the energy spectrum of the atomic chain with high precision through the photon scattering process. As an example, we demonstrate the efficacy of this method by simulating the Hofstadter butterfly spectrum at high resolution. Moreover, we find that for an incommensurate AAH atomic chain in the extended phase, the decays of the collective modes into the waveguide are concentrated on a few of these modes. In contrast, in the localized phase, the decays are distributed evenly in each mode. This interesting correspondence makes the localization transition easy to detect from the single-photon scattering spectrum. These results also indicate that wQED with giant atoms is optimal for investigating light-matter interactions between an AAH atomic chain and its surrounding photonic environment. The methodology presented here can be extended to simulate other types of 1D atomic chains.
@@
The remainder of this paper is organized as follows. In Sec. II, we derive the single-photon scattering amplitudes for wQED systems with multiple giant atoms based on the real-space formalism [65, 66]. Additionally, we extract the effective Hamiltonian that describes the interacting atomic chain. In Sec. III, we discuss quantum simulation of AAH model through wQED systems with giant atoms. Specifically, we present a method for constructing an AAH-type atomic chain based on giant-atom wQED structure in Sec. III.1. Subsequently, we provide discussions on probing the energy spectrum, simulating the Hofstadter butterfly, and probing the localization transition through single-photon scattering spectra in Sec. III.2-III.4. Additionally, we discuss the influence of spontaneous emission to the non-waveguide degrees of freedom in Sec. III.5. Finally, further discussions and conclusions are given in Sec. IV.
@@ II Theoretical description of single-photon scattering in waveguide QED containing multiple giant atoms
In this section, we provide an overview of the single-photon scattering problem for wQED with multiple giant atoms. From the derivation of the single-photon scattering amplitudes, we can extract the effective Hamiltonian describing the inter-atomic interactions mediated by the waveguide modes. This serves as the foundation for the construction of the target Hamiltonian that needs to be simulated. Additionally, in our proposal, single-photon scattering spectrum is also employed as a means to probe the many-body states of an atomic chain.
@@ II.1 Hamiltonian and equations of motion
Here we focus on the wQED structures comprising $N$ two-level giant atoms with modulated frequencies $\omega_{i}=\omega_{\text{a}}+V_{0}\cos\left(2\pi\beta{i}+\varphi\right)$ ($i=1,2,\cdots,N$). $V_{0}$ is the modulation amplitude, $\beta$ controls the periodicity of the modulation, and $\varphi$ is the modulation phase. Each atom is coupled to a 1D waveguide through two connection points and each pair of neighboring atoms is in a braided configuration, as shown in Fig. 1(a). The implementation of this structure can be achieved through a variety of schemes, such as coupling superconducting qubits with multiple interdigital transducers to surface acoustic waves in a waveguide via piezoelectric effects [5], or coupling superconducting qubits capacitively to a meandering microwave transmission line at multiple points [6]. And the frequency of each qubit can be modulated by a flux line. The connection points are located at position $x_{im}$ ($i=1,2,\cdots,N$, and $m=1,2$), the corresponding coupling strengths are $g_{i1}\equiv g_{1}$ and $g_{i2}\equiv g_{2}$, respectively. The phase delays between the coupling points are set as $\phi=\theta_{i2}-\theta_{i+1,1}$ and $\phi^{\prime}=\theta_{i+1,1}-\theta_{i1}$ ($\phi\leq\phi^{\prime}$), with $\theta_{im}=\omega_{\text{a}}x_{im}/v_{\text{g}}=k_{\text{a}}x_{im}$. $v_{\text{g}}$ is the group velocity of the photons in the waveguide. Under the rotating-wave approximation, the Hamiltonian of the system in the real space can be written as ($\hbar=1$)
$\hat{H}=$ $\sum_{i=1}^{N}{\omega_{i}\hat{\sigma}_{i}^{+}\!\>\hat{\sigma}_{i}^{-}}+\int\mathrm{d}x\sum_{s}{\hat{c}_{s}^{\dagger}(x)\left(-\mathrm{i}l_{s}v_{\text{g}}\frac{\partial}{\partial x}\right)\hat{c}_{s}\left(x\right)}$ (1)
[…]
where $s=\text{R},\text{L}$ and $l_{\text{R},\text{L}}=\pm 1$. $\hat{\sigma}_{i}^{+}$ ($\hat{\sigma}_{i}^{-}$) is the raising (lowering) operator of the atom $i$. $\hat{c}_{\text{R}}^{\dagger}(x)$ [$\hat{c}_{\text{R}}(x)$] and $\hat{c}_{\text{L}}^{\dagger}(x)$ [$\hat{c}_{\text{L}}(x)$] are the field operators of creating (annihilating) the right- and left-propagating photons at position $x$ in the waveguide.
@@
$|\Psi\rangle=\sum_{s}\int\mathrm{d}x\Phi_{s}(x)\hat{c}_{s}^{\dagger}(x)|\emptyset\rangle+\sum_{i=1}^{N}{f_{i}}\hat{\sigma}_{i}^{+}|\emptyset\rangle,$ (2)
@@
$\left(-\mathrm{i}v_{\text{g}}\frac{\partial}{\partial x}-\omega\right)\Phi_{\text{R}}(x)+\sum_{i=1}^{N}\sum_{m=1}^{2}g_{m}\delta\left(x-x_{im}\right)f_{i}=0,$ (4a)
$\left(\mathrm{i}v_{\text{g}}\frac{\partial}{\partial x}-\omega\right)\Phi_{\text{L}}(x)+\sum_{i=1}^{N}\sum_{m=1}^{2}g_{m}\delta\left(x-x_{im}\right)f_{i}=0,$ (4b)
@@ II.2 Scattering amplitudes of photons and effective Hamiltonian of atom array
We assume that a single photon with energy $\omega=v_{\text{g}}k$ is initially incident from the left, where $k$ is the wave vector of the photon. Then $\Phi_{\text{R}}(x)$ and $\Phi_{\text{L}}(x)$ take the following ansatz
$\Phi_{\text{R}}(x)=e^{\mathrm{i}kx}\sum_{p=0}^{2N}{t_{p}}\vartheta\left(x-x_{p}\right)\vartheta\left(x_{p+1}-x\right),$ (5a)
$\Phi_{\text{L}}(x)=e^{-\mathrm{i}kx}\sum_{p=1}^{2N}{r_{p}}\vartheta\left(x-x_{p-1}\right)\vartheta\left(x_{p}-x\right).$ (5b)
Here the positions of coupling points from left to right are labeled as $x_{1},x_{2},\cdots,x_{2N}$. Moreover, $x_{0}=-\infty$ and $x_{2N+1}=+\infty$ are defined. $t_{p}$ ($r_{p}$) is the transmission (reflection) amplitude of the $p$th coupling point. Particularly, $t_{2N}\equiv t$ ($r_{1}\equiv r$) is the transmission (reflection) amplitude of the last (first) coupling point. $t_{0}=1$ is the amplitude of the incident field. $\vartheta(x)$ denotes the Heaviside step function.
Starting from the equations of motion (4a)-(4c) and the ansatz (5a)-(5b), and after some algebra [32], we can obtain the transmission and reflection amplitudes
@@
The transmittance and the reflectance can be further defined as $T=|t|^{2}$ and $R=|r|^{2}$, respectively. In Eqs. (6a) and (6b), $\Delta=\omega-\omega_{\text{a}}$ is the photon-atom detuning, $\mathbf{I}$ is the identity matrix, and $\mathbf{V}$ takes the form
@@
Here the decay rate into the guided modes through the coupling point at $x_{im}$ is $\gamma_{m}=2g_{m}^{2}/v_{\text{g}}$. Note that we have assumed that the spacing between the connection points is small enough so that the phase-accumulated effects for detuned photons (the non-Markovian effects) can be neglected. Consequently, the wave vector $k$ in the definition of the phase factor has been replaced by $k_{\text{a}}$. $\mathbf{H}$ is the effective non-Hermitian Hamilton matrix of the atom array, with elements
$\mathcal{H}_{ij}=\left(\omega_{i}-\omega_{\text{a}}\right)\delta_{ij}-\frac{\mathrm{i}}{2}\sum_{m=1}^{2}\sum_{m^{\prime}=1}^{2}{{\sqrt{\gamma_{m}\gamma_{m^{\prime}}}}}e^{\mathrm{i}\left|\theta_{im}-\theta_{jm^{\prime}}\right|},$ (9)
which summarize the coherent and dissipative atom-atom interactions mediated by the waveguide modes. Moreover, this result indicates that the many-body model of the atomic chain can be designed by optimizing the layout of the connection points and the atom-waveguide coupling rates. Specifically, under the condition $\theta_{i2}-\theta_{i1}=\phi+\phi^{\prime}=(2n-1)\pi$ ($n\in\mathbb{Z}^{+}$), the Lamb shift of each atom vanishes, and the corresponding effective Hamiltonian of the atom array reads
$\hat{H}_{\text{eff}}$ $=$ $\sum_{i=1}^{N}\sum_{j=1}^{N}\mathcal{H}_{ij}\hat{\sigma}^{+}_{i}\hat{\sigma}^{-}_{j}$ (10)
$=$ $\sum_{i}\left[\left(\omega_{i}-\omega_{\text{a}}\right)-\frac{\mathrm{i}}{2}\Gamma_{\text{eff}}\right]\hat{\sigma}^{+}_{i}\hat{\sigma}^{-}_{i}$
$+\sum_{i\neq j}\left(G_{ij}-\frac{\mathrm{i}}{2}\Gamma^{(\text{coll})}_{ij}\right)\hat{\sigma}^{+}_{i}\hat{\sigma}^{-}_{j}.$
Here the effective decay of each atom is
$\Gamma_{\text{eff}}=2\left(\gamma-\sqrt{\gamma^{2}-\delta^{2}}\right),$ (11)
@@
$G_{ij}=\left\{\begin{array}[]{ll}\gamma{\sin}\phi\equiv{J},&|i-j|=1\\ \frac{(-1)^{|i-j|-1}}{2}\Gamma_{\text{eff}}\sin\left(|i-j|\phi\right),&|i-j|>1\end{array}\right.,$ (12a)
$\Gamma^{(\text{coll})}_{ij}=(-1)^{|i-j|-1}\Gamma_{\text{eff}}\cos[(i-j)\phi].$ (12b)
III Quantum simulation of AAH model through waveguide-QED systems with giant atoms
Figure 2: (a) Tansmitance for an AAH giant-atom chain with $N=30$ sites as a function of detuning $\Delta$ and modulation phase $\varphi$ for different values of $\beta$. From left to right, $\beta$ is chosen as $1/2$, $1/3$, $1/4$, and $(\sqrt{5}-1)/2$, respectively. The phase delay is selected to be $\phi=\pi/2$, resulting in $J=\gamma$. Other parameters are set as $\delta=0.1\gamma$ and $V_{0}=2J$. When $\beta$ is a rational number $\beta=p/q$ (with $p$ and $q$ being coprime), the corresponding spectrum is divided into $q$ bands (see the first three panels). On the other hand, when $\beta$ is irrational, the spectrum become fractal as a result of the quasiperiodicity of the system (see the last panel). (b) Energy spectra of an AAH model obtained from theoretical calculations for comparison with (a).
Figure 3: (a) Upper panel: transmission spectrum for an AAH giant-atom chain, with $\beta=1/4$ and $\varphi=3\pi/4$. Other parameters are the same as those used in Fig. 2 (a). Lower panel: the decomposed Lorentzian components for the spectrum shown in the upper panel. The resonances corresponding to the superradiant (subradiant) states in the band are labeled as SupS (SubS), with $\tilde{\Gamma}_{n}>\Gamma_{\text{eff}}$ ($\tilde{\Gamma}_{n}<\Gamma_{\text{eff}}$). The two edge states (ES) in the gaps exhibit a width that is approximately equal to $\Gamma_{\text{eff}}$. (b) The center frequency $\tilde{\Delta}_{n}$ (the disks) of each Lorentzian component and the eigenenergy $E_{n}$ (the crosses) given by the exact AAH model. (c) The linewidth $\tilde{\Gamma}_{n}$ of each Lorentzian component. The effective decay $\Gamma_{\text{eff}}$ of a single giant atom is labeled by the dashed line.